### v1

@@ III-C Estimation of the Effective Downlink Channel
A viable way to provide the receiver with CSI is to send downlink pilots that are precoded just as the data, so that user $k$ can estimate the effective channel $\mathbf{B}_{kk}$. Since the effective channel takes random non-Gaussian realizations from a stationary distribution, a suitable estimation method is LMMSE estimation [28, Ch. 12]. To enable the estimation, the APs jointly send orthonormal pilot sequences using the selected precoding. The pilot matrix assigned to user $k$ is $\boldsymbol{\Phi}_{k}\in\mathbb{C}^{\tau_{\mathrm{p}}\times M}$. Again, $\mathcal{P}_{k}$ is the set of users that use the same pilot matrix as user $k$. The pilot matrices are orthonormal, that is, $\boldsymbol{\Phi}_{i}^{\rm{H}}\boldsymbol{\Phi}_{i^{\prime}}=\mathbf{I}_{M}$ if $i,i^{\prime}\in\mathcal{P}_{k}$ and $\boldsymbol{\Phi}_{i}^{\rm{H}}\boldsymbol{\Phi}_{i^{\prime}}=\mathbf{0}$ otherwise. [^2] Note that we use the same pilot length $\tau_{\mathrm{p}}$ as in the uplink, but in principle, one could use a different length. By sending these pilots over the channel in (II-B1), the received signal at user $k$ becomes
@@
where $\mathbf{n}_{k}=\mathrm{vec}(\mathbf{N}_{k})$. We will let $\mathbf{b}_{kk}=\mathrm{vec}(\mathbf{B}_{kk})$ and $\mathbf{\tilde{y}}_{k}=\mathrm{vec}(\tilde{\mathbf{Y}}_{k})$ denote the vectorized effective channel and received signal, respectively. The LMMSE estimate of $\mathbf{b}_{kk}$ is [28, Ch. 12]
@@ III-D Capacity Bound with Estimated Channel at the Receiver
There are multiple ways that the user can select the receive combining matrix. One option is the MMSE combiner, which can be obtained by writing the MSE relation $\mathbb{E}\left\{|\mathbf{U}_{k}^{\rm{H}}\mathbf{y}_{k}-\boldsymbol{\varsigma}_{k}|^{2}\big{|}\hat{\mathbf{B}}_{ki}\;\mathrm{for}\;i=1,\ldots,K\right\}$ and finding the filter that minimizes it. The MMSE combiner for user $k$ can be written as
@@
The ZF combiner yields an easier-to-interpret SE expression than the MMSE combiner, but we anyway expect the latter to give larger SE values as it finds a better balance between signal equalization and interference suppression. A comparison between these different combining schemes will be given in Section VI.
@@ IV SE Results for Separate Stream Transmission
In this section, we will determine the SE relations for separate stream transmission. The relations have the same structure as in the former section, but can be rewritten to showcase the inner structure of the SE expressions. We note that the precoder structure must also change to reflect this change in SE expression. We will consider that in Section V, while the achievable SEs will be compared numerically in Section VI.
@@ IV-A Downlink Achievable SE without CSI at the Receiver
$\mathbf{y}_{k}=\bar{\mathbf{B}}^{\prime}_{kk}\boldsymbol{\Gamma}_{k}\boldsymbol{\varsigma}_{k}+\underbrace{(\mathbf{B}^{\prime}_{kk}-\bar{\mathbf{B}}^{\prime}_{kk})\boldsymbol{\Gamma}_{k}\boldsymbol{\varsigma}_{k}+\sum_{i=1,i\neq k}^{K}\mathbf{B}^{\prime}_{ki}\boldsymbol{\Gamma}_{i}\boldsymbol{\varsigma}_{i}+\mathbf{n}_{k}}_{\triangleq\mathbf{n}^{\prime}_{k}},$ (24)
@@ IV-C Downlink Achievable SE with Effective Channel Estimate at the Receiver
Next, we give an achievable SE relation with a general combiner for the same case as in the former section following a similar approach. The received signal in (II-B2) after combiner can be written as
$\tilde{\mathbf{y}}_{k}$ $=\underbrace{\mathbf{U}_{k}^{\rm{H}}\hat{\mathbf{B}}^{\prime}_{kk}}_{\triangleq\mathbf{E}^{\prime}_{kk}}\boldsymbol{\Gamma}_{k}\boldsymbol{\varsigma}_{k}+\mathbf{U}_{k}^{\rm{H}}\Big{(}\tilde{\mathbf{B}}^{\prime}_{kk}\boldsymbol{\Gamma}_{k}\boldsymbol{\varsigma}_{k}+\sum_{i=1,i\neq k}^{K}\mathbf{B}^{\prime}_{ki}\boldsymbol{\Gamma}_{i}\boldsymbol{\varsigma}_{i}+\mathbf{n}_{k}\Big{)}$
@@
$\hskip 70.0001pt+\mathbf{U}_{k}^{\rm{H}}\Big{(}\tilde{\mathbf{B}}^{\prime}_{kk}\boldsymbol{\Gamma}_{k}\boldsymbol{\varsigma}_{k}+\sum_{i=1,i\neq k}^{K}\mathbf{B}^{\prime}_{ki}\boldsymbol{\Gamma}_{i}\boldsymbol{\varsigma}_{i}+\mathbf{n}_{k}\Big{)}$
@@ V-B MMSE Precoding with CSI-Sharing and Separate Streams
$\overline{\mathbf{W}}^{\prime}_{lk}=q_{k}\left[\sum_{j=1}^{L}\sum_{i=1}^{K}q_{i}\Big{(}\hat{\mathbf{H}}_{ji}\boldsymbol{\Gamma}_{ji}\hat{\mathbf{H}}_{ji}^{\mathrm{H}}+\mathbf{C_{\tilde{\mathbf{H}}_{ji}\boldsymbol{\Gamma}_{ji}}}\Big{)}+\mathbf{I}_{N}\right]^{-1}\hat{\mathbf{H}}_{lk},$ (35)
@@
The derivation of this precoder can be found in Appendix A-B. The precoder $\mathbf{W}_{k}$ obtained with this transmission method is block diagonal as the precoders $\mathbf{W}_{lk}=\mathbf{W}^{\prime}_{lk}\boldsymbol{\Gamma}_{lk}$ found here are stacked to form a block diagonal matrix. The $\boldsymbol{\Gamma}_{ji}$ matrices select the specific columns of each channel to which transmission is allowed and suppress interference in those directions. Moreover, we only consider the channel estimation errors in those directions.
@@ V-D A Channel Condition for Equivalent Performance of Same Stream and Separate Stream Transmissions
where $\mathbf{U}_{k}=[\mathbf{U}_{1k},\ldots,\mathbf{U}_{L_{k}k}]$, $\boldsymbol{\Sigma}_{k}=\rm{diag}(\boldsymbol{\Sigma}_{1k},\ldots,\boldsymbol{\Sigma}_{L_{k}k})$ and $\mathbf{V}_{k}=\rm{diag}(\mathbf{V}_{1k},\ldots,\mathbf{V}_{L_{k}k})$. We assume $L_{k}=L=M$ is the total number of APs serving user $k$ for this example case. For a channel with this structure, the matrix $\mathbf{V}_{k}^{\mathrm{H}}$ with the right singular vectors is block diagonal with each block associated with one AP. This implies that signals transmitted from two different APs cannot be received in the same subspace and therefore cannot add constructively. Consequently, when the channel has the block-diagonal structure in (38), the optimal same stream transmission precoder automatically becomes block diagonal and coincides with the optimal separate stream transmission precoder. A channel with this structure can appear exactly or approximately when the user sees the APs in very different angular directions, for example, in a connected stadium where a few APs are equally spaced around the perimeter of the stadium.
VI Numerical Results
In this section, we provide numerical results where we will compare the downlink SEs achieved with different transmission methods and different numbers of antennas and APs. We use the same simulation setup as in [4]. Hence, we assume that the AP and user locations are uniformly distributed in a $1\times 1\;\text{km}^{2}$ area. We use $\tau_{\mathrm{p}}$ mutually orthogonal pilot sequences that are shared among the users. Pilot sharing is done as explained in [4, Ch. 4] and each pilot sequence is used by $K^{\prime}$ users, which is an integer value chosen as a ratio of $K$. Therefore, the pilot length is $\tau_{\mathrm{p}}=K^{\prime}M$. The same pilot length and assignment are used for uplink and downlink channel estimation. For downlink transmission, the power control coefficients explained in Section V are used.
@@ VI Numerical Results
Table I summarizes the main differences between these transmission methods, which are related to what information must be shared between the APs when implementing them.
TABLE I: Comparison of Options
@@
CSI Sharing $\boldsymbol{\times}$
Data Sharing $\boldsymbol{\times}$$\boldsymbol{\times}$
Moreover, the total number of streams $M$ is less than the number of APs $L$. We assume that $L=20$, $N=4$, $M=2$, and $K=5$ unless stated otherwise. Furthermore, we assume that $\tau_{p}=KM/2$ and the correlated Rayleigh fading channels are modeled as in [4, Sec. 5.2] with an angular standard deviation (ASD) of $15^{\circ}$. We consider the bounds we derived with the downlink channel estimate for the results unless otherwise stated. The total number of AP antennas bounds the multiplexing gain. With $M$ antennas per user and $K$ users, our configuration supports up to $MK$ data streams, which is the metric to compare with previous works. Our setup can be viewed as a subset of a larger cell-free massive MIMO network operating on a specific time-frequency resource block. In practice, resource allocation and scheduling partition a larger number of APs and UEs into smaller groups, ensuring that not all users are active simultaneously.
VI-A Differences Between the Proposed Methods
We first investigate the impact of the choice of transmission methods on the SEs. Fig. 2 shows the average SE of a randomly located user as a function of the number of antennas per AP with the MMSE precoders. From this figure, we can see that Method 1 performs better than the other methods. This outcome is anticipated as Method 1 builds on sharing CSI among all APs so that the centralized MMSE precoder can be utilized to suppress interference in a coordinated manner. Method 3 provides slightly lower SEs than Method 1, as the lack of data sharing forces the precoding matrix to be block diagonal. However, Method 3 can still suppress inter-AP interference since CSI is shared to enable that. One clear benefit of Method 3 is that there is no need for phase synchronization between the APs, while for Method 1, synchronization is crucial for the APs to jointly beamforming signals that add constructively at the intended user and destructively at other users. Finally, Method 2 performs worse than Method 3 as it lacks CSI sharing between APs, but it can only suppress inter-AP interference based on its statistics.
@@ VI-A Differences Between the Proposed Methods
VI-B Impact of Spatial Correlation
Next, we investigate the impact spatial correlation has on the SEs. Spatial correlation is modeled similarly to that in [4, Sec. 5.2] assuming that both AP and users are equipped with ULAs. Moreover, the spatial correlation matrix elements are computed between each AP and user antennas as explained there. The ASD variation changes the correlation from the AP’s viewpoint, thereby modeling how widely the scattering objects are distributed around the user. Fig. 3 shows the average SE as a function of the ASD, where a small value means high spatial correlation and vice versa. The figure shows that when the ASD is low, the same stream transmission and separate stream transmission with CSI sharing work particularly well. The performance then reduces as the ASD increases. However, the separate stream transmission without CSI sharing (Method 2) performs better when the ASD is higher. Hence, spatial correlation can either enhance or degrade the system’s performance depending on the method used.

### v2

@@ III-C Estimation of the Effective Downlink Channel
A viable way to provide the receiver with CSI is to send downlink pilots that are precoded just as the data, so that user $k$ can estimate the effective channel $\mathbf{B}_{kk}$. Since the effective channel takes random non-Gaussian realizations from a stationary distribution, a suitable estimation method is LMMSE estimation [kay1993fundamentals, Ch. 12]. To enable the estimation, the APs jointly send orthonormal pilot sequences using the selected precoding. The pilot matrix assigned to user $k$ is $\boldsymbol{\Phi}_{k}\in\mathbb{C}^{\tau_{\mathrm{p}}\times M}$. Again, $\mathcal{P}_{k}$ is the set of users that use the same pilot matrix as user $k$. The pilot matrices are orthonormal, that is, $\boldsymbol{\Phi}_{i}^{\rm{H}}\boldsymbol{\Phi}_{i^{\prime}}=\mathbf{I}_{M}$ if $i,i^{\prime}\in\mathcal{P}_{k}$ and $\boldsymbol{\Phi}_{i}^{\rm{H}}\boldsymbol{\Phi}_{i^{\prime}}=\mathbf{0}$ otherwise. [^2] Note that we use the same pilot length $\tau_{\mathrm{p}}$ as in the uplink, but in principle, one could use a different length. By sending these pilots over the channel in (II-B1), the received signal at user $k$ becomes
@@
where $\mathbf{n}_{k}=\mathrm{vec}(\mathbf{N}_{k})$. We will let $\mathbf{b}_{kk}=\mathrm{vec}(\mathbf{B}_{kk})$ and $\mathbf{\tilde{y}}_{k}=\mathrm{vec}(\tilde{\mathbf{Y}}_{k})$ denote the vectorized effective channel and received signal, respectively. The LMMSE estimate of $\mathbf{b}_{kk}$ is [kay1993fundamentals, Ch. 12]
@@ III-D Capacity Bound with Estimated Channel at the Receiver
We note that the SE relation captures both the data rate during data transmission and the loss from the pilot overhead. Resulting effects of this and the improvement in SE will be shown in Section VII. There are multiple ways that the user can select the receive combining matrix. One option is the MMSE combiner, which can be obtained by writing the MSE relation $\mathbb{E}\left\{|\mathbf{U}_{k}^{\rm{H}}\mathbf{y}_{k}-\boldsymbol{\varsigma}_{k}|^{2}\big|\hat{\mathbf{B}}_{ki}\;\mathrm{for}\;i=1,\ldots,K\right\}$ and finding the filter that minimizes it. The MMSE combiner for user $k$ can be written as
@@
The ZF combiner yields an easier-to-interpret SE expression than the MMSE combiner, but we anyway expect the latter to give larger SE values as it finds a better balance between signal equalization and interference suppression. A comparison between these different combining schemes will be given in Section VII.
@@ IV SE Results for Separate Stream Transmission
In this section, we will determine the SE relations for separate stream transmission. The relations have the same structure as in the former section, but can be rewritten to showcase the inner structure of the SE expressions. We note that the precoder structure must also change to reflect this change in SE expression. We will consider that in Section V, while the achievable SEs will be compared numerically in Section VII.
@@ IV-A Downlink Achievable SE without CSI at the Receiver
$\mathbf{y}_{k}=\bar{\mathbf{B}}^{\prime}_{kk}\boldsymbol{\Gamma}_{k}\boldsymbol{\varsigma}_{k}+\underbrace{(\mathbf{B}^{\prime}_{kk}-\bar{\mathbf{B}}^{\prime}_{kk})\boldsymbol{\Gamma}_{k}\boldsymbol{\varsigma}_{k}+\hskip-5.0pt\sum_{i=1,i\neq k}^{K}\mathbf{B}^{\prime}_{ki}\boldsymbol{\Gamma}_{i}\boldsymbol{\varsigma}_{i}+\mathbf{n}_{k}}_{\triangleq\mathbf{n}^{\prime}_{k}},$ (24)
@@ IV-C Downlink Achievable SE with Effective Channel Estimate at the Receiver
Next, we give an achievable SE relation with a general combiner for the same case as in the former section, following a similar approach. The received signal in (II-B2) after combiner can be written as
$\tilde{\mathbf{y}}_{k}$ $=\underbrace{\mathbf{U}_{k}^{\rm{H}}\hat{\mathbf{B}}^{\prime}_{kk}}_{\triangleq\mathbf{E}^{\prime}_{kk}}\boldsymbol{\Gamma}_{k}\boldsymbol{\varsigma}_{k}+\mathbf{U}_{k}^{\rm{H}}\Big(\tilde{\mathbf{B}}^{\prime}_{kk}\boldsymbol{\Gamma}_{k}\boldsymbol{\varsigma}_{k}+\hskip-5.0pt\sum_{i=1,i\neq k}^{K}\hskip-5.0pt\mathbf{B}^{\prime}_{ki}\boldsymbol{\Gamma}_{i}\boldsymbol{\varsigma}_{i}+\mathbf{n}_{k}\Big)$
@@
$\hskip 70.0001pt+\mathbf{U}_{k}^{\rm{H}}\Big(\tilde{\mathbf{B}}^{\prime}_{kk}\boldsymbol{\Gamma}_{k}\boldsymbol{\varsigma}_{k}+\hskip-5.0pt\sum_{i=1,i\neq k}^{K}\hskip-5.0pt\mathbf{B}^{\prime}_{ki}\boldsymbol{\Gamma}_{i}\boldsymbol{\varsigma}_{i}+\mathbf{n}_{k}\Big)$
@@ V-B MMSE Precoding with CSI-Sharing and Separate Streams
$\overline{\mathbf{W}}^{\prime}_{lk}=q_{k}\left[\sum_{j=1}^{L}\sum_{i=1}^{K}q_{i}\Big(\hat{\mathbf{H}}_{ji}\boldsymbol{\Gamma}_{ji}\hat{\mathbf{H}}_{ji}^{\mathrm{H}}+\mathbf{C_{\tilde{\mathbf{H}}_{ji}\boldsymbol{\Gamma}_{ji}}}\Big)+\mathbf{I}_{N}\right]^{-1}\hskip-5.0pt\hat{\mathbf{H}}_{lk},$ (35)
@@
The derivation of this precoder can be found in Appendix A-B. The precoder $\mathbf{W}_{k}$ obtained with this transmission method is block diagonal, as the precoders $\mathbf{W}_{lk}=\mathbf{W}^{\prime}_{lk}\boldsymbol{\Gamma}_{lk}$ found here are stacked to form a block diagonal matrix. The $\boldsymbol{\Gamma}_{ji}$ matrices select the specific columns of each channel to which transmission is allowed and suppress interference in those directions. Moreover, we only consider the channel estimation errors in those directions.
@@ V-D A Channel Condition for Equivalent Performance of Same Stream and Separate Stream Transmissions
where $\mathbf{U}_{k}=[\mathbf{U}_{1k},\ldots,\mathbf{U}_{L_{k}k}]$, $\boldsymbol{\Sigma}_{k}=\rm{diag}(\boldsymbol{\Sigma}_{1k},\ldots,\boldsymbol{\Sigma}_{L_{k}k})$ and $\mathbf{V}_{k}=\rm{diag}(\mathbf{V}_{1k},\ldots,\mathbf{V}_{L_{k}k})$. We assume $L_{k}=L=M$ is the total number of APs serving user $k$ for this example case. For a channel with this structure, the matrix $\mathbf{V}_{k}^{\mathrm{H}}$ with the right singular vectors is block diagonal, with each block associated with one AP. This implies that signals transmitted from two different APs cannot be received in the same subspace and therefore cannot add constructively. Consequently, when the channel has the block-diagonal structure in (38), the optimal same stream transmission precoder automatically becomes block diagonal and coincides with the optimal separate stream transmission precoder. A channel with this structure can appear exactly or approximately when the user sees the APs in very different angular directions, for example, in a connected stadium where a few APs are equally spaced around the perimeter of the stadium.
VI Computational Costs
We will now quantify the computational complexity per coherence block when using the proposed schemes. The matrices used in uplink channel estimation depend on the statistics of the channels. Processing the received signal during the pilot transmission requires $N\tau_{p}$ complex multiplication per pilot sequence, and the MMSE channel estimation process requires $(NM)^{2}$ complex multiplications. Therefore, the complexity for the uplink channel estimation is $K((NM)^{2}+N\tau_{p})$ complex multiplications per coherence block.
The complexity of computing the precoder matrices is dominated by the $LN\times LN$ matrix inversions. The number of complex multiplications per coherence block is
$\frac{(NL)^{2}+NL}{2}MK+(NL)^{2}M+\frac{(NL)^{3}-NL}{3}.$
From this relation, we can see that the complexity increases linearly with $\tau_{p}$ as it is proportional to $M$
TABLE I: comparison of fronthaul communication load (total number of complex scalars) between the same-stream and the separate-stream transmission schemes.
Scheme Pilot Signaling Data Signaling
Same-stream $\tau_{p}NM$ $(\tau_{c}-\tau_{p})N$
Separate-stream $\tau_{p}NM$ $(\tau_{c}-\tau_{p})N$
The fronthaul signaling related to sending the received pilot signals from each AP to the CPU is $\tau_{p}N$ complex scalars for the same stream transmission case. The fronthaul signaling for the precoded downlink data signals necessitates sending $N\times 1$ vectors comprising the multiplication of the precoder matrix and the data vector. Therefore, the fronthaul load for the data signals is $(\tau_{c}-\tau_{p})N$ complex scalars. This value is irrespective of the choice of the precoding scheme.
For the separate stream case, the fronthaul signaling for the uplink channel estimation and sending the data is the same. This happens as we multiply the precoder matrix by the data vector. A comparison of these can be seen in Table I.
VII Numerical Results
In this section, we provide numerical results where we will compare the downlink SEs achieved with different transmission methods and different numbers of antennas and APs. We use the same simulation setup as in [demir2021foundations]. Hence, we assume that the AP and user locations are uniformly distributed in a $1\times 1\;\text{km}^{2}$ area. We use $\tau_{\mathrm{p}}$ mutually orthogonal pilot sequences that are shared among the users. Pilot assignment is done as explained in [demir2021foundations, Ch. 4] and each pilot sequence is used by $K^{\prime}$ users, which is an integer value chosen as a ratio of $K$. Therefore, the pilot length is $\tau_{\mathrm{p}}=K^{\prime}M$. The same pilot length and assignment are used for uplink and downlink channel estimation. We adopt the pilot assignment scheme from [demir2021foundations], as simple allocation methods are remarkably effective in cell-free systems. This efficacy stems from the pronounced pathloss variations inherent to the architecture, which create a natural spatial separation among users and thereby intrinsically limit pilot contamination. Furthermore, our use of MMSE channel estimation provides a powerful, built-in mechanism for decontamination that uses the channel covariance matrices. As established in [8094949], the MMSE estimator statistically separates interfering users by exploiting their distinct spatial channel statistics. While our approach already exploits these statistics for decontamination, one could alternatively use the covariance matrices $\mathbf{R}_{lk}$ for pilot design in cell-free massive MIMO systems. For instance, a correlation-aware pilot assignment method could reduce pilot overhead or contamination by assigning pilots based on the similarity of users’ spatial correlation structures, ensuring that users with overlapping channel subspaces receive minimally interfering pilots. This mitigates inter-user interference during estimation without necessitating additional pilot resources. Such approaches are explored in [9160993, 6415397]. Moreover, subspace-based training methods could reduce pilot overhead or contamination by projecting pilots onto the dominant eigenspaces of $\mathbf{R}_{lk}$, allowing for training sequences that target the low-rank channel structure rather than the full dimension, which is particularly effective for multi-antenna users to lower the required pilot length while improving estimation accuracy in the presence of spatial correlation. These techniques are demonstrated in [9110802, 11044434]. We do not incorporate these additional techniques in this study in order to maintain a manageable system complexity and, more importantly, to clearly isolate the performance gains attributable to the core contributions of this paper. For downlink transmission, the power control coefficients explained in Section V are used.
@@ VII Numerical Results
Table II summarizes the main differences between these transmission methods, which are related to what information must be shared between the APs when implementing them.
TABLE II: Comparison of Options
@@
[…]
Moreover, the total number of streams $M$ is less than the number of APs $L$. We assume that $L=20$, $N=4$, $M=2$, and $K=5$ unless stated otherwise. Furthermore, we assume that $\tau_{p}=KM/2$ and the correlated Rayleigh fading channels are modeled as in [demir2021foundations, Sec. 5.2] with an angular standard deviation (ASD) of $15^{\circ}$. We consider the bounds we derived with the downlink channel estimate for the results, unless otherwise stated. The total number of AP antennas bounds the multiplexing gain. With $M$ antennas per user and $K$ users, our configuration supports up to $MK$ data streams, which is the metric to compare with previous works. Our setup can be viewed as a subset of a larger cell-free massive MIMO network operating on a specific time-frequency resource block. In practice, resource allocation and scheduling partition a larger number of APs and UEs into smaller groups, ensuring that not all users are active simultaneously.
VII-A Differences Between the Proposed Methods
We first investigate the impact of the choice of transmission methods on the SEs. Fig. 2 shows the average SE of a randomly located user as a function of the number of antennas per AP with the MMSE precoders. From this figure, we can see that Method 1 performs better than the other methods. This outcome is anticipated as Method 1 builds on sharing CSI among all APs so that the centralized MMSE precoder can be utilized to suppress interference in a coordinated manner. Method 3 provides slightly lower SEs than Method 1, as the lack of data sharing forces the precoding matrix to be block diagonal. However, Method 3 can still suppress inter-AP interference since CSI is shared to enable that. One clear benefit of Method 3 is that there is no need for phase synchronization between the APs, while for Method 1, synchronization is crucial for the APs to jointly beamform signals that add constructively at the intended user and destructively at other users. Finally, Method 2 performs worse than Method 3 as it lacks CSI sharing between APs, but it can only suppress inter-AP interference based on its statistics.
@@ VII-A Differences Between the Proposed Methods
VII-B Impact of Spatial Correlation
Next, we investigate the impact spatial correlation has on the SEs. Spatial correlation is modeled similarly to that in [demir2021foundations, Sec. 5.2] assuming that both AP and users are equipped with ULAs. Moreover, the spatial correlation matrix elements are computed between each AP and user antennas as explained there. The ASD variation changes the correlation from the AP’s viewpoint, thereby modeling how widely the scattering objects are distributed around the user. Fig. 3 shows the average SE as a function of the ASD, where a small value means high spatial correlation and vice versa. The figure shows that when the ASD is low, the same stream transmission and separate stream transmission with CSI sharing work particularly well. The performance then reduces as the ASD increases. However, the separate stream transmission without CSI sharing (Method 2) performs better when the ASD is higher. Hence, spatial correlation can either enhance or degrade the system’s performance depending on the method used.